# Multi-objective Aerial IRS-assisted ISAC Optimization via Generative AI-enhanced Deep Reinforcement Learning

### v1

@@ Abstract
Integrated sensing and communication (ISAC) has garnered substantial research interest owing to its pivotal role in advancing the development of next-generation (6G) wireless networks. However, achieving a performance balance between communication and sensing in the dual-function radar communication (DFRC)-based ISAC system remains a significant challenge. In this paper, an aerial intelligent reflecting surface (IRS)-assisted ISAC system is explored, where a base station (BS) supports dual-functional operations, enabling both data transmission for multiple users and sensing for a blocked target, with the channel quality enhanced by an IRS mounted on the unmanned aerial vehicle (UAV). Moreover, we formulate an integrated communication, sensing, and energy efficiency multi-objective optimization problem (CSEMOP), which aims to maximize the communication rate of the users and the echo rate of the target, while minimizing UAV propulsion energy consumption by jointly optimizing the BS beamforming matrix, IRS phase shifts, the flight velocity and angle of the UAV. Considering the non-convexity, trade-off, and dynamic nature of the formulated CSEMOP, we propose a generative diffusion model-based deep deterministic policy gradient (GDMDDPG) method to solve the problem. Specifically, the diffusion model is incorporated into the actor network of DDPG to improve the action quality, with noise perturbation mechanism for better exploration and recent prioritized experience replay (RPER) sampling mechanism for enhanced training efficiency. Simulation results indicate that the GDMDDPG method delivers superior performance compared to the existing methods.
@@
I Introduction
As fifth-generation (5G) networks commercially deploys and research on sixth-generation (6G) progresses, conventional communication systems have revealed notable shortcomings in meeting the increasingly diverse and growing demands [1, 2]. In particularly, as the low-altitude economy continues to evolve [3], leveraging airspace for activities such as logistics and rescue operations, the need for seamless communication and real-time sensing becomes even more critical. Therefore, it is necessary to adopt a new approach that moves beyond traditional communication-centric network designs and embraces integrated sensing-communication networks. In this case, integrated sensing and communication (ISAC) technology stands out as a practical and efficient solution. [4], and it has been recognized by the International Telecommunication Union (ITU) as a critical application scenario among the six primary use cases envisioned for the 6G era [5].
In the practical implement of ISAC systems, non-ideal propagation environments caused by multipath fading or obstacle occlusion can significantly affect system performance [6]. In this context, intelligent reflecting surface (IRS) emerges as a promising solution [7, 8]. Specifically, IRS consists of a planar array with densely packed sub-wavelength reflective elements, each of which can individually modify the phase and amplitude of incoming signals. [9]. Because each reflective element is small in size, an IRS with practical size can incorporate numerous elements to enable substantial beamforming gains via co-modulation, effectively compensating for path loss or constructing virtual line-of sight (LoS) links [10]. Thus, the unique capability for remodeling propagation environment provides an innovative solution to enhance the robustness of the ISAC system in occlusion scenarios [11]. Moreover, forming an aerial IRS by integrating the IRS on a maneuverable unmanned aerial vehicle (UAV) can break through the spatial coverage limitation of the traditional fixed IRS, so that further enhancing the effectiveness of the ISAC system [12].
@@ I Introduction
However, designing an effective beamforming and resource allocation strategy address the competing demands between the communication and sensing functionalities in DRFC architecture remains challenging. [16]. In particular, in the aerial IRS-assisted ISAC system, dynamic characteristics are introduced to both application scenario and optimization problem due to the time-varying nature of channels and the mobility of aerial IRS. However, obtaining accurate prior knowledge in advance in the dynamic scenario is difficult, which makes traditional optimization methods (e.g., convex optimization) show limitations in solving such dynamic optimization problem [17]. In this context, deep reinforcement learning (DRL) provides a feasible solution. Specifically, DRL enables the agent to learn by interacting with the environment, removing the necessity for prior information and attaining performance close to optimal. However, traditional DRL algorithms employ Multi-Layer Perceptron (MLP)-based actor networks, which struggle to capture the complex environment features of the aerial IRS-assisted ISAC system [18]. Moreover, DRL algorithms typically necessitate a substantial number of agent-environment interactions to achieve meaningful learning progress.
@@
To address the above challenges, we first formulate a multi-objective optimization problem, and propose a generative artificial intelligence (GenAI)-enabled DRL method. To the best of our knowledge, no existing work simultaneously optimizes communication, sensing, and energy efficiency performance in the movable IRS-assisted ISAC system. The primary contributions of this paper are outlined as follows:
@@
  • Aerial IRS-assisted ISAC System: An aerial IRS-assisted ISAC system is explored, where a dual-functional base station (BS) equipped with multiple antennas simultaneously communicates with multiple users and provides sensing services to an occluded target via beamforming. Moreover, to mitigate the impact of obstacles on the sensing performance of the BS, the IRS is mounted on a UAV to form an aerial IRS. Compared with fixed IRS installations, aerial IRS offers greater spatial flexibility, enabling the dynamic establishment of virtual LoS links in response to environment changes, thereby significantly improving the channel quality of the system.
@@
  • Formulation of Multi-objective Optimization Problem: An integrated communication, sensing, and energy efficiency multi-objective optimization problem (CSEMOP) is formulated, which aims at maximizing the communication rate of users and the echo rate of the target while minimizing the UAV propulsion energy consumption by jointly optimizing the BS active beamforming matrix, IRS phase shifts, the flight velocity and angle of the UAV. Notably, the formulated CSEMOP is a non-convex dynamic optimization problem with trade-offs, which makes traditional optimization algorithms limited in solving the problem.
@@
  • GenAI-enabled DRL Method: We first transform the formulated CSEMOP into a Markov decision process (MDP), and then propose a diffusion model-based deep deterministic policy gradient (GDMDDPG) method, which is a GenAI-enabled DRL method to solve the problem. Specifically, the diffusion model-based actor network integrates the analytical and generative capabilities of the diffusion model, enhancing both environment state analysis and action generation of the method. Moreover, the noise perturbation mechanism introduces stochasticity to overcome policy stagnation and foster exploration within the action space. Meanwhile, the RPER mechanism combines recent experience emphasis (ERE) and prioritized experience replay (PER) to dynamically select high-value recent experiences during network updates, which accelerates learning process of the method.
@@
  • Simulation and Analysis: Simulation results validate the effectiveness of the proposed GenAI-enabled DRL method. Specifically, the proposed method exhibits superior performance across various simulation settings compared to the three deployment methods and two other DRL benchmark algorithms, particularly achieving enhanced communication and sensing performance even under lower UAV energy consumption. Moreover, we perform a simulation analysis of the proposed method across various UAV initial locations to further validate its robustness.
@@
The organization of this paper is as follows: Section II introduces related work. Section III presents the system model. Section IV formulates and analyzes the optimization problem. Section V provides the proposed GenAI-enabled DRL method. Section VI conducts the simulations. Finally, the paper is summarized in Section VII.
II Related Work
@@
II-A IRS-assisted ISAC System
@@
For IRS-assisted DFRC-based system, the authors in [19] utilized semi-passive IRS to assist single-antenna users in achieving single-input multiple-output (SIMO) millimeter-wave (mmWave) communication with a BS while simultaneously performing location-aware services. The authors in [20] deployed active IRS equipped with amplifiers to assist multiple-input multiple-output (MIMO) DFRC-based ISAC scenario. Moreover, multi-IRS-assisted ISAC systems are explored in [21], addressing two distinct sensing scenarios involving point targets and extended targets. In addition, the authors in [22] explored the scenario where sensing signals enable communication with multiple users, taking into account the influence of environmental clutter on ISAC systems supported by IRS. However, the aforementioned existing works primarily employ fixed IRS to assist ISAC system. In contrast, mobile IRS can provide more flexible support for ISAC system, which can further enhances the system performance.
II-B Optimization Variables in IRS-assisted ISAC System
@@
In terms of communication metrics, the authors in [23, 12] optimized resource allocation for the IRS-supported DFRC-based ISAC system to improve the communication rate for users. Moreover, the authors in [24, 25] tackled eavesdropping threats in the ISAC scenario by refining BS beamforming matrix and adjusting IRS phase shifts to enhance the secure communication rate. In terms of sensing metrics, the authors in [16] explored serving users and targets simultaneously by transmitting superimposed NOMA signals, aiming to maximize the beampattern gain, which measures energy concentration on the target. Based on this, the authors in [26] further explored maximizing the beampattern gain for targets in a STAR-RIS-supported ISAC system with environmental clutter. Moreover, the authors in [27] explored ISAC technology to tackle precise vehicle positioning and high-quality communication challenges by optimizing IRS phase shifts to minimize the position error bound. However, the aforementioned existing works optimized either communication or sensing metrics individually, without addressing the simultaneous optimization of both metrics.
II-C Optimization Methods in IRS-assisted ISAC System
@@
Currently, various optimization methods have been widely used to solve the optimization problems in IRS-assisted ISAC systems. For example, the authors in [28] adopted a multi-strategy alternating optimization algorithm based on quadratically constrained quadratic programming and semidefinite relaxation to enhance both the transmission rate and probing power. Moreover, the authors in [29] designed an improved particle swarm optimization algorithm to jointly optimize active and passive beamforming for maximizing the echo signal power. In addition, the authors in [30] employed the proximal policy optimization (PPO) algorithm to optimize beamforming matrix and the UAV trajectory to maximize the transmission rate in the aerial IRS-assisted ISAC system. Furthermore, the authors in [31] explored a STAR-RIS-supported secure ISAC framework and used the soft actor-critic (SAC) and DDPG algorithms to improve the long-term secrecy rate.
@@
However, both convex optimization and swarm intelligence methods face limitations in dynamic scenarios where the prior knowledge is difficult to acquire. Moreover, while DRL algorithms show inherent advantages over traditional optimization methods, limitations still persist in analyzing complex environmental features and state-action relationships, which may affect the quality of decision.
Figure 1: Aerial IRS-assisted ISAC system.
III System Model
@@
In this section, we provide detailed models of communication, sensing and UAV mobility in the aerial IRS-assisted ISAC system.
III-A System Overview
@@
As shown in Fig. 1, an aerial IRS-assisted ISAC system is considered, which consists of a BS equipped with a uniform antenna array of $M$ antennas, a ground user set denoted as $\mathcal{N}=\{1,\cdots,N\}$, and a ground target. Specifically, the dual-functional BS communicates with the $N$ users while performing sensing tasks for the target. We assume that the presence of obstacles prevents direct sensing link between the BS and the target, which significantly affects the sensing performance of the BS [12]. In this case, we deploy a UAV-carried IRS, consisting of $L$ reflective elements denoted as =$\mathcal{L}=\{1,\cdots,L\}$, to establish virtual line-of-sight (LoS) links between the BS and the target, which can mitigate the blocking effect of existing obstacles. Therefore, the introduction of the aerial IRS enhances both sensing and communication performance by improving signal quality, which boosts a more flexible and robust system.
@@
We discretize the operation time of the system, dividing the total time $T_{\text{d}}$ into $T$ equal-length time slots, denoted as $\mathcal{T}=\{1,\cdots,T\}$, with the length of each time slot $t_{\text{d}}=\frac{T_{\text{d}}}{T}$. Moreover, we adopt the widely used Cartesian coordinate system to describe the locations of elements in our scenario. Specifically, due to the maneuverability of the UAV, the location of the aerial IRS is variable, and it can be represented as $q^{\text{r}}[t]=(x^{\text{r}}[t],y^{\text{r}}[t],z^{\text{r}})$ in time slot $t$. Similarly, the locations of the BS, $n$-th user and the target are denoted as $q^{\text{b}}=(x^{\text{b}},y^{\text{b}},z^{b})$, $q_{n}^{\text{u}}=(x_{n}^{\text{u}},y_{n}^{\text{u}},0)$, $q^{\text{g}}=(x^{\text{g}},y^{\text{g}},0)$, respectively.
@@
Notably, multiple sensing and communication links change as the trajectory of the aerial IRS varies, which introduces highly dynamic to the considered system. In the following, we provide a detailed model for sensing, communication, and UAV mobility to capture the key optimization variables and objectives.
III-B Channel Model
@@
The channel between the BS and the user is comprised of two main components, i.e., the direct link and the reflected link assisted by the aerial IRS. Next, a detailed description for the channel model is provided.
III-B1 Direct Link
@@
Due to the intricate nature of the signal propagation in the considered scenario, we utilize the Rician model to represent the direct links, effectively capturing both the LoS component and scattering effects from multipath propagation. [12, 32]. Then, the channel vector $h_{n}^{bu}[t]\in\mathbb{C}^{M\times 1}$ from the BS to $n$-th user in time slot $t$ can be represented as
@@
where $L_{0}$ represents the channel power gain at a reference distance of 1 m, $\alpha^{\text{bu}}$ denotes the path loss, and $\eta^{\text{bu}}$ is the Rician factor. Moreover, $d_{n}^{\text{bu}}=\sqrt{\|q^{\text{b}}-q_{n}^{\text{u}}\|^{2}}$ represents the distance between the BS and the $n$-th user. In addition, $g_{n,\text{L}}^{\text{bu}}[t]$ denotes the LoS component, and $g_{n,\text{N}}^{\text{bu}}[t]\sim\mathcal{N}(0,\mathbf{I})$ is the NLoS component that follows the complex Gaussian distribution.
III-B2 Reflected Link
Note that the reflected link consists of the BS-aerial IRS link and the aerial IRS-users link, respectively.
When the aerial IRS is deployed at a sufficiently high altitude, the associated air-ground communication links are primarily dominated by the LoS component [32]. Then, the channel vector $h^{\text{br}}[t]\in\mathbb{C}^{L\times M}$ between the BS and the aerial IRS in time slot $t$ is given by

### v2

# Multi-objective Low-altitude IRS-assisted ISAC Optimization via Generative AI-enhanced Deep Reinforcement Learning

@@ Abstract
Integrated sensing and communication (ISAC) has garnered substantial research interest owing to its pivotal role in advancing the development of next-generation (6G) wireless networks. However, achieving a performance balance between communication and sensing in the dual-function radar communication (DFRC)-based ISAC system remains a significant challenge. In this paper, a low-altitude intelligent reflecting surface (IRS)-assisted ISAC system is explored, where a base station (BS) supports dual-functional operations, enabling both data transmission for multiple users and sensing for a blocked target, with the channel quality enhanced by an IRS mounted on the unmanned aerial vehicle (UAV). Moreover, we formulate an integrated communication, sensing, and energy efficiency multi-objective optimization problem (CSEMOP), which aims to maximize the communication rate of the users and the sensing rate of the target, while minimizing UAV propulsion energy consumption by jointly optimizing the BS beamforming matrix, IRS phase shifts, the flight velocity and angle of the UAV. Considering the non-convexity, trade-off, and dynamic nature of the formulated CSEMOP, we propose a generative diffusion model-based deep deterministic policy gradient (GDMDDPG) algorithm to solve the problem. Specifically, the diffusion model is incorporated into the actor network of DDPG to improve the action quality, with noise perturbation mechanism for better exploration and recent prioritized experience replay (RPER) sampling mechanism for enhanced training efficiency. Simulation results indicate that the GDMDDPG algorithm delivers superior performance compared to the existing methods.
@@
1 Introduction
As fifth-generation (5G) networks commercially deploys and research on sixth-generation (6G) progresses, conventional communication systems have revealed notable shortcomings in meeting the increasingly diverse and growing demands [1, 2]. In particular, as the low-altitude economy continues to evolve [3], leveraging airspace for activities such as logistics and rescue operations, the need for seamless communication and real-time sensing becomes even more critical. Therefore, it is necessary to adopt a new approach that moves beyond traditional communication-centric network designs and embraces integrated sensing-communication networks. In this case, integrated sensing and communication (ISAC) technology stands out as a practical and efficient solution [4], and it has been recognized by the International Telecommunication Union (ITU) as a critical application scenario among the six primary use cases envisioned for the 6G era [5].
In the practical implementation of ISAC systems, non-ideal propagation environments caused by multipath fading or obstacle occlusion can significantly affect system performance [6]. In this case, the intelligent reflecting surface (IRS) emerges as a promising solution [7, 8]. Specifically, IRS consists of a planar array with densely packed sub-wavelength reflective elements, each of which can individually modify the phase and amplitude of incoming signals [9, 10]. Because each reflective element is small in size, an IRS with practical size can incorporate numerous elements to enable substantial beamforming gains via co-modulation, effectively compensating for path loss or constructing virtual line-of-sight (LoS) links [11]. Thus, the unique capability for remodeling propagation environment provides an innovative solution to enhance the robustness of the ISAC system in occlusion scenarios [12]. Moreover, forming a low-altitude IRS by integrating the IRS on a maneuverable unmanned aerial vehicle (UAV) can overcome the spatial coverage limitation of the traditional fixed IRS, so that further enhancing the effectiveness of the ISAC system [13].
@@ 1 Introduction
However, designing an effective beamforming and resource allocation strategy addressing the competing demands between the communication and sensing functionalities in DRFC architecture remains challenging [17]. In particular, in the low-altitude IRS-assisted ISAC system, dynamic characteristics are introduced to both application scenario and optimization problem due to the time-varying nature of channels and the mobility of low-altitude IRS. However, obtaining accurate prior knowledge in advance in the dynamic scenario is difficult, which makes traditional optimization methods (e.g., convex optimization) show limitations in solving such dynamic optimization problem [18]. In this context, deep reinforcement learning (DRL) provides a promising solution. Specifically, DRL enables the agent to learn by interacting with the environment, removing the necessity for prior information and attaining performance close to optimal [19]. However, traditional DRL algorithms employ Multi-Layer Perceptron (MLP)-based actor networks, which struggle to capture the complex environment characteristics of the low-altitude IRS-assisted ISAC system [20, 21]. Moreover, DRL algorithms typically necessitate a substantial number of agent-environment interactions to achieve meaningful learning progress.
@@
To address the above challenges, we first formulate a multi-objective optimization problem, and propose a generative artificial intelligence (GenAI)-enabled DRL algorithm. To the best of our knowledge, no existing work simultaneously optimizes communication, sensing, and energy efficiency performance in the movable low-altitude IRS-assisted ISAC system. The main contributions of this paper are outlined as follows:
@@
  • Low-altitude IRS-assisted ISAC System: A low-altitude IRS-assisted ISAC system is explored, where a dual-functional base station (BS) equipped with multiple antennas simultaneously communicates with multiple users and provides sensing services to an occluded target via beamforming. Moreover, to mitigate the impact of obstacles on the sensing performance of the BS, the IRS is mounted on a UAV to form a low-altitude IRS. Compared with fixed IRS installations, low-altitude IRS offers greater spatial flexibility, enabling the dynamic establishment of virtual LoS links in response to environment changes, thereby significantly improving the channel quality of the system.
@@
  • Formulation of Multi-objective Optimization Problem: An integrated communication, sensing, and energy efficiency multi-objective optimization problem (CSEMOP) is formulated, which aims at maximizing the communication rate of users and the sensing rate of the target while minimizing the UAV propulsion energy consumption by jointly optimizing the BS active beamforming matrix, IRS phase shifts, the flight velocity and angle of the UAV. Notably, the formulated CSEMOP is a non-convex dynamic optimization problem, which makes traditional optimization algorithms limited in solving the problem.
@@
  • GenAI-enabled DRL Algorithm: We first transform the formulated CSEMOP into a Markov decision process (MDP), and then propose a diffusion model-based deep deterministic policy gradient (GDMDDPG) algorithm, which is a GenAI-enabled DRL algorithm to solve the problem. Specifically, the diffusion model-based actor network integrates the analytical and generative capabilities of the diffusion model, enhancing both environment state analysis and action generation of the algorithm. Moreover, the noise perturbation mechanism introduces stochasticity to overcome policy stagnation and foster exploration within the action space. Meanwhile, the RPER mechanism combines recent experience emphasis (ERE) and prioritized experience replay (PER) to dynamically select high-value recent experiences during network updates, which accelerates learning process of the algorithm.
@@
  • Simulation and Analysis: Simulation results validate the effectiveness of the proposed GenAI-enabled DRL algorithm. Specifically, the proposed algorithm exhibits superior performance across various simulation settings compared to the three deployment methods and two other DRL benchmark algorithms, particularly achieving enhanced communication and sensing performance even under lower UAV energy consumption. Moreover, we perform a simulation analysis of the proposed algorithm across various UAV initial locations to further validate its robustness.
@@
The organization of this paper is as follows: Section 2 introduces the related work. Section 3 presents the system model. Section 4 formulates and analyzes the optimization problem. Section 5 provides the proposed GenAI-enabled DRL algorithm. Section 6 conducts the simulations. Finally, the paper is summarized in Section 7.
2 Related Work
@@
2.1 IRS-assisted ISAC System
@@
For IRS-assisted DFRC-based system, the authors in [22] utilized semi-passive IRS to assist single-antenna users in achieving single-input multiple-output (SIMO) millimeter-wave (mmWave) communication with a BS while simultaneously performing location-aware services. The authors in [23] deployed active IRS equipped with amplifiers to assist multiple-input multiple-output (MIMO) DFRC-based ISAC scenario. Moreover, multi-IRS-assisted ISAC systems are explored in [24], addressing two distinct sensing scenarios involving point targets and extended targets. In addition, the authors in [25] explored the scenario where sensing signals enable communication with multiple users, taking into account the influence of environmental clutter on ISAC systems supported by IRS.
However, the aforementioned existing works primarily employ fixed IRS to assist ISAC system. In contrast, mobile IRS can provide more flexible support for the ISAC system, which can further enhance the system performance.
2.2 Optimization Variables in IRS-assisted ISAC System
@@
In terms of the communication metrics, the authors in [26, 13] optimized resource allocation for the IRS-supported DFRC-based ISAC system to improve the communication rate for users. Moreover, the authors in [27, 28] tackled eavesdropping threats in the ISAC scenario by refining BS beamforming matrix and adjusting IRS phase shifts to enhance the secure communication rate. For the sensing metrics, the authors in [17] explored serving users and targets simultaneously by transmitting superimposed NOMA signals, aiming to maximize the beampattern gain, which measures energy concentration on the target. Based on this, the authors in [29] further explored maximizing the beampattern gain for targets in a STAR-RIS-supported ISAC system with environmental clutter. Moreover, the authors in [30] explored ISAC technology to tackle vehicle positioning and high-quality communication challenges by optimizing IRS phase shifts to minimize the position error bound.
However, the above existing works optimized either communication or sensing metrics individually, without addressing the simultaneous optimization of both metrics.
2.3 Optimization Methods in IRS-assisted ISAC System
@@
Various optimization methods have been widely used to solve the optimization problems in IRS-assisted ISAC systems. For example, the authors in [31] adopted a multi-strategy alternating optimization algorithm based on quadratically constrained quadratic programming and semidefinite relaxation to enhance both the transmission rate and probing power. Moreover, the authors in [32] designed an improved particle swarm optimization algorithm to jointly optimize active and passive beamforming for maximizing the echo signal power. In addition, the authors in [33] employed the proximal policy optimization (PPO) algorithm to optimize the beamforming matrix and the UAV trajectory to maximize the transmission rate in the low-altitude IRS-assisted ISAC system. Furthermore, the authors in [34] explored a STAR-RIS-supported secure ISAC framework and used the soft actor-critic (SAC) and DDPG algorithms to improve the long-term secrecy rate.
@@
However, both convex optimization and swarm intelligence methods face limitations in dynamic scenarios where the prior knowledge is difficult to acquire [18]. Moreover, while DRL algorithms show inherent advantages over traditional optimization methods, limitations still persist in analyzing complex environmental characteristics and state-action relationships, which may affect the quality of decision.
Figure 1: Low-altitude IRS-assisted ISAC system.
3 System Model
@@
In this section, we provide detailed models of communication, sensing, and UAV mobility in the low-altitude IRS-assisted ISAC system.
3.1 System Overview
@@
As shown in Fig. 1, a low-altitude IRS-assisted ISAC system is considered, which consists of a BS equipped with a uniform antenna array of $M$ antennas, a ground user set denoted as $\mathcal{N}=\{1,\ldots,N\}$, and a ground target. Specifically, the dual-functional BS communicates with $N$ users while performing sensing tasks for the target. We assume that obstacles prevent the direct sensing link between the BS and the target, which significantly affects the sensing performance of the BS [13]. In this case, we deploy a UAV-carried IRS as a low-altitude IRS, consisting of $L$ reflective elements denoted as =$\mathcal{L}=\{1,\ldots,L\}$, to establish virtual line-of-sight (LoS) links between the BS and the target, which can mitigate the blocking effect of existing obstacles. Therefore, the introduction of the low-altitude IRS enhances both sensing and communication performance by improving signal quality, which boosts a more flexible and robust system.
@@
We discretize the operation time of the system, dividing the total time $T_{\text{d}}$ into $T$ equal-length time slots, denoted as $\mathcal{T}=\{1,\cdots,T\}$, with the length of each time slot $t_{\text{d}}=\frac{T_{\text{d}}}{T}$. Moreover, we adopt the widely used Cartesian coordinate system to describe the locations of elements in our scenario. Specifically, due to the maneuverability of the UAV, the location of the low-altitude IRS is variable, and it can be represented as $q^{\text{r}}[t]=(x^{\text{r}}[t],y^{\text{r}}[t],z^{\text{r}})$ in time slot $t$. Similarly, the locations of the BS, $n$-th user and the target are denoted as $q^{\text{b}}=(x^{\text{b}},y^{\text{b}},z^{b})$, $q_{n}^{\text{u}}=(x_{n}^{\text{u}},y_{n}^{\text{u}},0)$, $q^{\text{g}}=(x^{\text{g}},y^{\text{g}},0)$, respectively.
@@
Notably, multiple sensing and communication links change as the UAV trajectory varies, which introduces highly dynamic to the considered system. In the following, we provide a detailed model for sensing, communication, and UAV mobility to capture the key optimization variables and objectives.
3.2 Channel Model
@@
The channel between the BS and the user is comprised of two main components, i.e., the direct link and the reflected link assisted by the low-altitude IRS. Next, a detailed description for the channel model is provided.
3.2.1 Direct Link
@@
Due to the intricate nature of the signal propagation in the considered scenario, we utilize the Rician model to represent the direct links, effectively capturing both the LoS component and scattering effects from multipath propagation [13, 35]. Then, the channel vector $h_{n}^{bu}[t]\in\mathbb{C}^{M\times 1}$ from the BS to $n$-th user in time slot $t$ can be represented as
@@
where $L_{0}$ represents the channel power gain at a reference distance of 1 m, $\alpha^{\text{bu}}$ denotes the path loss, and $\eta^{\text{bu}}$ is the Rician factor. Moreover, $d_{n}^{\text{bu}}=\sqrt{\|q^{\text{b}}-q_{n}^{\text{u}}\|^{2}}$ represents the distance between the BS and the $n$-th user. In addition, $g_{n,\text{L}}^{\text{bu}}[t]$ denotes the LoS component, and $g_{n,\text{N}}^{\text{bu}}[t]\sim\mathcal{CN}(0,\mathbf{I})$ is the NLoS component that follows the complex Gaussian distribution, and $\mathbf{I}$ is the identity matrix.
3.2.2 Reflected Link
Note that the reflected link consists of the BS-low-altitude IRS link and the low-altitude IRS-users link, respectively.
When the low-altitude IRS is deployed at a sufficiently high altitude, the associated air-ground communication links are primarily dominated by the LoS component [35]. Then, the channel vector $h^{\text{br}}[t]\in\mathbb{C}^{L\times M}$ between the BS and the low-altitude IRS in time slot $t$ is given by